3.) Westerly Industries has the following financial information. What is its cash conversion​ cycle? (Hint: Use a​ 365-day year.)
Sales   
​$140, 000
Cost of Goods Sold
112,000
Accounts Receivable
56,000
Inventory
22, 400
Accounts Payable
56,000
The cash conversion cycle is ____ days.  ​(Round to two decimal​ places.)
Answer #1

Inventory days = (22,400/112,000)365 = 73

Payable days = (56,000/112,000)365 = 182.50

Receivable days = (56,000/140,000)365 = 146

Cash conversion days = 73 + 146 - 182.50

Cahs Conversion Cycle = 36.50 days